# TELEClass: Taxonomy Enrichment and LLM-Enhanced Hierarchical Text Classification with Minimal Supervision

### v1

@@ 3.3. Core Class Refinement with Enriched Taxonomy
Unlike previous methods in flat text classification (Wang et al., 2021a) that use keyword-level embeddings to estimate document and class representations, here, because we already have a set of roughly classified documents for each class, $D^{0}_{c}$, we can directly define their representations based on document-level embeddings. To obtain document representations, we utilize a pretrained Sentence Transformer model (Reimers and Gurevych, 2019) to encode the entire document, which we denote as $\vec{d}$. Then, for each class $c$, we identify a subset of its assigned documents which explicitly mention at least one of the class-indicative keywords and thus most confidently belong to this class, $D_{c}=\{d\in D^{0}_{c}|\exists w\in T_{c},w\in d\}$. Then, we use the average of their document embeddings as the class representation, $\vec{c}=\frac{1}{|D_{c}|}\sum_{d\in D_{c}}\vec{d}$. Finally, we compute the document-class matching score as the cosine similarity between their representations.
@@
(6) $\small\begin{split}conf_{i}=\text{diff}^{\,i}(m_{i}),\quad\mathbb{C}_{i}&=\{c^{i}_{1},c^{i}_{2},\dots,c^{i}_{m_{i}}\},\\ m_{i}=\operatorname*{arg\,max}_{j\in\{1,\dots,|\mathcal{C}|-1\}}&\text{diff}^{\,i}(j).\end{split}$
@@
Finally, we select the set of documents and their refined core classes with the top 50% confidence scores $conf_{i}$, to be the final set of refined core classes for classifier training. Formally, $\mathcal{D}^{\text{core}}=\{d_{i}\}$, where each $d_{i}\in\mathcal{D}^{\text{core}}$ satisfies |$|\{d_{j}|conf_{j}<conf_{i}\}|\geq\frac{1}{2}|\mathcal{D}|$.
@@
3.4. Text Classifier Training with Path-Based Data Augmentation
@@
The final step of TELEClass is to train a hierarchical text classifier using the confident refined core classes. One straightforward way is to directly use the selected core classes as a complete set of pseudo-labeled documents and train a text classifier in a common supervised way. However, such a strategy is ineffective, because the core classes are not comprehensive enough and cannot cover all the classes in the taxonomy. The reason is two-fold. First, the core class selection step is document-centric, meaning that it only focuses on selecting the most accurate core classes for each document, while ignoring whether each class in the taxonomy is ever selected as a core class. Second, the label space in the hierarchical text classification task naturally contains some fine-grained and long-tail classes, so they are often not guaranteed to be selected as core classes due to their low frequency. Empirically, for the two datasets we use in our experiments, Amazon and DBPedia, the missing rates of the selected core classes are 11.6% and 5.4%, respectively. These missing classes will never be used as positive classes in the training process, if we only train the classifier with the selected core classes.
@@
Therefore, to overcome this issue, we further utilize the language generation ability of LLMs to augment the current pseudo-labels. We propose the idea of path-based document generation by LLMs to generate augmented documents for each path from a level-1 node to a leaf node in the taxonomy. By adding the generated documents to the pseudo-labeled data, we can ensure the coverage of the pseudo-data. Moreover, we use a path instead of a single class to guide the LLM generation, because the meaning of lower-level classes is often conditioned on their parents. For example, in Figure 2, a path “hair care” $\rightarrow$ “shampoo” can guide the LLM to generate text about hair shampoo instead of pet shampoo or carpet shampoo that are in different paths.
More specifically, for each path from a level-1 node to a leaf node in the taxonomy, $p=(c_{1},\dots,c_{l})$, we query an LLM to generate a small number of documents (e.g., $q=5$) corresponding to the path $p$ and denote the resulting documents as $d^{p}_{i}$, $i=1,\dots,q$. By generating $q$ documents for each path and treating all the classes in the query path as the pseudo-labels for the generated documents, we get a set of generated data $\mathcal{D}^{\text{gen}}=\{d^{p}_{i}\}_{p\in Path(\mathcal{T}),i=1,\dots,q}$, where $Path(\mathcal{T})$ denotes all such paths in the taxonomy $\mathcal{T}$, and $\mathbb{C}_{p}$ is the set of classes in the path $p$. $\mathcal{D}^{\text{gen}}$ can serve as augmented data to train the text classifier together with the selected refined core classes $\mathcal{D}^{\text{core}}$. Notice that, here we generate pseudo documents for each path instead of just generating for the classes that have no or very few assigned documents. The reason is that, because our core classes are obtained based on static keyword-based features (i.e., the enriched class-indicative topical terms), it may induce certain bias of the matching process into the pseudo-data and affect the learning process (Dong et al., 2023). Therefore, we use an LLM to generate pseudo documents for each path in the taxonomy, which can also serve as regularization of the potential bias.
@@
Now, with two sets of pseudo-data prepared, we are ready to introduce the classifier architecture and the training process.
Classifier architecture. Because we do not focus on proposing a new model for the hierarchical text classification task, we choose to use a standard architecture as our text classifier, including a document encoder, class embeddings, and a matching network.
  • Document encoder: we initialize the document encoder $g(\cdot)$ with a pretrained BERT-base model (Devlin et al., 2019) and use the hidden representation of the [CLS] token from the last layer as the document encoding, $\mathbf{d}_{i}=g(d_{i})$.
  • Class embeddings: we initialize the class embeddings with the class surface name embeddings obtained by a pretrained BERT model (c.f. Section 3.2), denoted by $\mathbf{c}_{j}$. Notice that, here we detach the embeddings from the base model, so only the embeddings will be updated during the training process without back-propagation to the backbone model.
  • Matching Network: we utilize a log-bilinear matching network to calculate the probability of document $d_{i}$ belonging to class $c_{j}$:
    $\small p(c_{j}|d_{i})=\mathcal{P}(y_{j}=1|d_{i})=\sigma(\exp{(\mathbf{c}_{j}^{T}\mathbf{W}\mathbf{d}_{i})}),$
    where $\sigma$ stands for the sigmoid function and $\mathbf{W}$ is a learnable interaction matrix.
@@
Training process. For each document $d_{i}$ tagged with core classes $\mathbb{C}_{i}$, we construct its positive classes $\mathbb{C}^{\text{core}}_{i,+}$ as the union of its core classes and their ancestors in the label taxonomy, and its negative classes $\mathbb{C}^{\text{core}}_{i,-}$ are the ones that are not positive classes or descendants of any core class. This is because the ancestors of confident core classes are also likely to be true labels, and the descendants may not all be negative given that the automatically generated core classes are not optimal. Formally,
$\small\mathbb{C}^{\text{core}}_{i,+}$ $=\mathbb{C}_{i}\cup\left(\cup_{c\in\mathbb{C}_{i}}Anc(c)\right),$
$\mathbb{C}^{\text{core}}_{i,-}$ $=\mathcal{C}-\mathbb{C}^{\text{core}}_{i,+}-\cup_{c\in\mathbb{C}_{i}}Des(c),$
where $Anc(c)$ and $Des(c)$ denote the set of ancestors and descendants and class $c$, respectively. For the LLM-generated documents, we are more confident in its pseudo labels, so we simply treat all the classes in the corresponding path as positive classes and all other classes as negative,
$\small\mathbb{C}^{\text{gen}}_{p,+}=\mathbb{C}_{p},\quad\mathbb{C}^{\text{gen}}_{p,-}=\mathcal{C}-\mathbb{C}_{p}.$
Then, we train a multi-label classifier with the binary cross entropy loss (BCE):
$\mathcal{L}^{\text{core}}=-\sum_{d_{i}\in\mathcal{D}^{\text{core}}}\left(\sum_{c_{j}\in\mathbb{C}^{\text{core}}_{i,+}}\log{p(c_{j}|d_{i})}+\sum_{c_{j}\in\mathbb{C}^{\text{core}}_{i,-}}\log{(1-p(c_{j}|d_{i}))}\right)$
$\mathcal{L}^{\text{gen}}=-\sum_{d^{p}_{i}\in\mathcal{D}^{\text{gen}}}\left(\sum_{c_{j}\in\mathbb{C}^{\text{gen}}_{p,+}}\log{p(c_{j}|d^{p}_{i})}+\sum_{c_{j}\in\mathbb{C}^{\text{gen}}_{p,-}}\log{(1-p(c_{j}|d^{p}_{i}))}\right)$
(7) $\small\mathcal{L}=\mathcal{L}^{\text{core}}+\frac{|\mathcal{D}^{\text{core}}|}{|\mathcal{D}^{\text{gen}}|}\cdot\mathcal{L}^{\text{gen}},$
where $\frac{|\mathcal{D}^{\text{core}}|}{|\mathcal{D}^{\text{gen}}|}$ serves as a scaling factor to balance two sets of pseudo-data and can be exactly calculated as $\frac{|\mathcal{D}|}{2q\cdot|Path(\mathcal{T})|}$.
@@
We train the text-matching model with $\mathcal{L}$ to get the final classifier. Notice that, we do not continue to train the classifier with self-training that is commonly used in previous studies (Meng et al., 2019; Shen et al., 2021), because we want to keep the classifier training part simple. Using self-training may further improve the model performance, which we leave for future exploration. Algorithm 1 summarizes TELEClass.

### v2

@@ 3.3 Core Class Refinement with Enriched Taxonomy
To obtain document representations, we utilize a pre-trained Sentence Transformer model Reimers and Gurevych (2019) to encode the entire document, which we denote as $\vec{d}$. Then, for each class $c$, we identify a subset of its assigned documents that explicitly mention at least one of the class-indicative keywords and thus most confidently belong to this class, $D_{c}=\{d\in D^{0}_{c}|\exists w\in T_{c},w\in d\}$. Then, we use the average of their document embeddings as the class representation, $\vec{c}=\frac{1}{|D_{c}|}\sum_{d\in D_{c}}\vec{d}$. Finally, we compute the document-class matching score as the cosine similarity between their representations.
@@
$\begin{split}conf_{i}=\text{diff}^{\,i}(m_{i}),\quad\mathbb{C}_{i}&=\{c^{i}_{1},\dots,c^{i}_{m_{i}}\},\\ m_{i}=\operatorname*{arg\,max}_{j\in\{1,\dots,|\mathcal{C}|-1\}}&\text{diff}^{\,i}(j).\end{split}$ (2)
@@
Finally, we select the top 75% of documents $d_{i}$ and their refined core classes with the highest confidence scores $conf_{i}$, denoted as $\mathcal{D}^{\text{core}}$.
@@
3.4 Text Classifier Training with Path-Based Data Augmentation
@@
The final step of TELEClass is to train a hierarchical text classifier using the confident refined core classes. One straightforward way is to directly use the selected core classes as a complete set of pseudo-labeled documents and train a text classifier in a common supervised way. However, such a strategy is ineffective, because the core classes are not comprehensive enough and cannot cover all the classes in the taxonomy. This is because the hierarchical label space naturally contains fine-grained and long-tail classes, and they are often not guaranteed to be selected as core classes due to their low frequency. Empirically, for the two datasets we use in our experiments, Amazon and DBPedia, the percentages of classes never selected as core classes are 11.6% and 5.4%, respectively. These missing classes will never be used as positive classes in the training process if we only train the classifier with the selected core classes.
@@
Therefore, to overcome this issue, we propose the idea of path-based document generation by LLMs to generate a small number of augmented documents (e.g., $q=5$) for each distinct path from a level-1 node to a leaf node in the taxonomy. By adding the generated documents to the pseudo-labeled data, we can ensure that each class of the taxonomy will be a positive class of at least $q$ documents. Moreover, we use a path instead of a single class to guide the LLM generation, because the meaning of lower-level classes is often conditioned on their parents. For example, in Figure 2, a path “hair care” $\rightarrow$ “shampoo” can guide the LLM to generate text about hair shampoo instead of pet shampoo or carpet shampoo that are in different paths. To promote data diversity, we make one LLM query for each path and ask it to generate $q$ diverse documents. Appx. C shows the prompts we used. We denote the generated documents as $\mathcal{D}^{\text{gen}}$.
@@
Now, with two sets of data, the pseudo-labeled documents $\mathcal{D}^{\text{core}}$ and LLM-generated documents $\mathcal{D}^{\text{gen}}$, we are ready to introduce the classifier architecture and the training process.
Classifier architecture. We use a simple text matching network Shen et al. (2021) as our model architecture, which includes a document encoder initialized with a pre-trained BERT-base model Devlin et al. (2019) and a log-bilinear matching network. Class representations are initialized by class name embeddings (c.f. Sec. 3.2).
@@
Training process. For each document with refined core classes, we construct its positive classes as the union of its refined core classes and their ancestors in the label taxonomy, and its negative classes are the ones that are not positive classes or descendants of any core class. This is because the ancestors of confident core classes are also likely to be true labels, and the descendants may not all be negative given that the automatically generated core classes are not optimal. For the LLM-generated documents, we are more confident in their labels, so we simply treat all the classes in the corresponding label path as positive classes and all other classes as negative.
@@
We train the text-matching model with the standard binary cross-entropy loss. The loss terms of two sets of data are weighted by their relative size, $\frac{|\mathcal{D}^{\text{core}}|}{|\mathcal{D}^{\text{gen}}|}$. Notice that we do not continue training the classifier with self-training that is commonly used in previous studies Meng et al. (2019); Shen et al. (2021). Using self-training may further improve the model performance, which we leave for future exploration. Algorithm 1 in Appendix summarizes TELEClass.
Table 1: Experiment results on Amazon-531 and DBPedia-298 datasets, evaluated by Example-F1, P@k, and MRR. The best score among zero-shot and weakly-supervised methods is boldfaced. “” indicates the numbers for these baselines are directly from previous paper Shen et al. (2021). “—” means the method cannot generate a ranking of predictions and thus MRR cannot be calculated.
Supervision Type Methods Amazon-531 DBPedia-298
Example-F1 P@1 P@3 MRR Example-F1 P@1 P@3 MRR
Zero-Shot Hier-0Shot-TC 0.4742 0.7144 0.4610 0.6765 0.7871 0.6765
ChatGPT 0.5164 0.6807 0.4752 0.4816 0.5328 0.4547
Weakly-Supervised Hier-doc2vec 0.3157 0.5805 0.3115 0.1443 0.2635 0.1443
WeSHClass 0.2458 0.5773 0.2517 0.3047 0.5359 0.3048
TaxoClass-NoST 0.5431 0.7918 0.5414 0.5911 0.7712 0.8621 0.7712 0.8221
TaxoClass 0.5934 0.8120 0.5894 0.6332 0.8156 0.8942 0.8156 0.8762
TELEClass 0.6483 0.8505 0.6421 0.6865 0.8633 0.9351 0.8633 0.8864
4 Experiments
4.1 Experiment Setup
Datasets. We use two public datasets in different domains for evaluation. (1) Amazon-531 McAuley and Leskovec (2013) consists of Amazon product reviews and a three-layer label taxonomy of 531 product types, with 29,487 unlabeled training samples and 19,685 testing samples. (2) DBPedia-298 Lehmann et al. (2015) consists of Wikipedia articles with a three-layer label taxonomy of 298 categories, with 196,665 unlabeled training samples and 49,167 testing samples.
Compared Methods. We compare TELEClass with the following methods: Hier-0Shot-TC Yin et al. (2019), zero-shot ChatGPT (i.e., GPT-3.5-turbo), Hier-doc2vec Le and Mikolov (2014), WeSHClass Meng et al. (2019), TaxoClass Shen et al. (2021) and its variant TaxoClass-NoST that does not apply self-training. We also include a Fully-Supervised baseline that uses the entire labeled training data to train the text-matching network. See more details of these methods in Appex. B.1. The implementation details of TELEClass are in Appx. B.3.
Evaluation Metrics. Following previous studies Shen et al. (2021), we utilize the following evaluation metrics: Example-F1, Precision at k (P@k) for $k=1,3$, and Mean Reciprocal Rank (MRR). See Appendix B.2 for detailed definitions.
4.2 Experimental Results
Table 1 shows the evaluation results of all the compared methods. We make the following observations. (1) Overall, TELEClass achieves significantly better performance than other strong zero-shot and weakly-supervised baselines. (2) By comparing with other weakly-supervised methods, we find that TELEClass significantly outperforms TaxoClass-ST, the strongest baseline that, like TELEClass, does not use self-training. This further demonstrates the effectiveness of TELEClass. (3) Although LLMs (e.g., ChatGPT) show superior power in many tasks, naïvely prompting it in the hierarchical text classification task yields significantly inferior performance compared to strong weakly-supervised text classifiers. We conduct a more detailed comparison with LLM prompting for hierarchical text classification in Section 4.4.
Table 2: Performance of TELEClass and its ablations on Amazon-531 and DBPedia-298.
Methods Amazon-531 DBPedia-298
Example-F1 P@1 P@3 MRR Example-F1 P@1 P@3 MRR
Gen-Only 0.5151 0.7477 0.5096 0.5357 0.7930 0.9421 0.7930 0.8209
TELEClass-NoLLMEnrich 0.5520 0.7370 0.5463 0.5900 0.8319 0.9108 0.8319 0.8563
TELEClass-NoCorpusEnrich 0.6143 0.8358 0.6082 0.6522 0.8185 0.8916 0.8185 0.8463
TELEClass-NoGen 0.6449 0.8348 0.6387 0.6792 0.8494 0.9187 0.8494 0.8730
TELEClass 0.6483 0.8505 0.6421 0.6865 0.8633 0.9351 0.8633 0.8864
Table 3: Performance comparison of TELEClass and zero-shot LLM prompting.
Methods Amazon-531 DBPedia-298
F1 P@1 P@3 F1 P@1 P@3
ChatGPT 0.516 0.681 0.475 0.482 0.533 0.455
ChatGPT-L 0.662 0.857 0.644 0.665 0.830 0.649
GPT-4 0.699 0.822 0.689 0.605 0.652 0.592
TELEClass 0.648 0.851 0.642 0.863 0.935 0.863
4.3 Ablation Studies
We conduct ablation studies to better understand how each component of TELEClass contributes to final performance. Table 2 shows the results of the following ablations:
  • Gen-Only only uses the augmented documents by path-based LLM generation to train the final classifier.
  • TELEClass-NoLLMEnrich excludes the LLM-based taxonomy enrichment component.
  • TELEClass-NoCorpusEnrich excludes the corpus-based taxonomy enrichment component.
  • TELEClass-NoGen excludes the augmented documents by path-based LLM generation.
We find that the full model TELEClass achieves the overall best performance among the compared methods, showing the effectiveness of each of its components. First, both the LLM-based and corpus-based enrichment modules bring improvement to the performance. Interestingly, we find that they make different levels of contribution on the two datasets: LLM-based enrichment brings more improvement on Amazon-531 while corpus-based enrichment contributes more on DBPedia-298. We suspect the reasons are as follows. The classes in Amazon-531 are commonly seen product types that LLM can understand and enrich in a reliable manner. However, DBPedia-298 contains classes that are more subtle to distinguish, which can also be shown by the lower performance of zero-shot LLM prompting on DBPedia compared to Amazon-531 (c.f. ChatGPT in Table 1). Therefore, corpus-based enrichment can consolidate the meaning of each class based on corpus-specific knowledge to facilitate better classification. We also find that path-based LLM generation consistently improves the model performance while requiring only a few hundred queries to LLMs. Even Gen-Only achieves comparable performance to the strong baseline TaxoClass-NoST, demonstrating the effectiveness of this augmentation step.
4.4 Comparison with Zero-Shot LLM Prompting
In this section, we further compare TELEClass with zero-shot LLM prompting. Because it is not straightforward to get ranked predictions by LLMs, we only report Example-F1 and P@k as performance evaluations. We report the zero-shot performance of ChatGPT and GPT-4 by including all the classes in the prompt and asking the LLM to provide the three most appropriate classes for a given document. Experiments on GPT-4 are conducted on 1000 randomly sampled test documents given the prohibitively high cost. We also include a level-by-level version, ChatGPT-L, which selects one most appropriate class at each level and continues to its children.
Table 3 shows the experiment results. We find that TELEClass consistently outperforms zero shot performance of LLMs by large margins on DBPedia-298. On Amazon-531, while it continues to show substantial improvements over ChatGPT, it slightly underperforms ChatGPT-L and GPT-4. As for the cost, once trained, TELEClass does not require an additional cost on inference and also has a substantially shorter inference time. Prompting LLMs takes a longer time and can be prohibitively expensive, and the cost will scale up with an increasing size of test data. The estimated API cost for ChatGPT on the entire DBPedia test set is \$80 and for GPT-4 it is around \$2,500. Also, we find that the level-by-level version ChatGPT-L consistently outperforms the naïve version, demonstrating the necessity of taxonomy structure.
Table 4: Intermediate results on two documents, including selected core classes by different methods, true labels, and the corresponding taxonomy enrichment results. The optimal core class in the true labels is marked with ©.
Dataset Document Core Classes by… True Labels Corr. Enrichment
DBPedia The Lindenhurst Memorial Library (LML) is located in Lindenhurst, New York, and is one of the fifty six libraries that are part of the Suffolk Cooperative Library System … TaxoClass: village TELEClass initial: building TELEClass refined: library library©, agent, educational institution Class: library Enrichment: national library, central library, collection, volumes…
Amazon Since mom (89 yrs young) isn’t steady on her feet, we have place these grab bars around the room. It gives her the stability and security she needs. TaxoClass: personal care, health personal care, safety TELEClass initial: daily living aids, medical supplies equipment, safety, TELEClass refined: bathroom aids safety health personal care, medical supplies equipment, bathroom aids safety© Class: bathroom aids safety Enrichment: seat, toilet, shower, safety, handles…
4.5 Case studies
To better understand the TELEClass framework, we show some intermediate results of two documents in Table 4, including the core classes selected by (1) the original TaxoClass Shen et al. (2021) method, (2) TELEClass’s initial core classes selected by LLM, and (3) refined core classes of TELEClass. Besides, we also include the true labels of the documents and the taxonomy enrichment results of the corresponding core class in the table. Overall, we can see that TELEClass’s refined core class is the most accurate. For example, for the first Wikipedia article about a library, TaxoClass selects “village” as the core class, while TELEClass’s initial core class finds a closer one “building” thanks to the power of LLMs. Then, with the enriched classes-indicative features as guidance, TELEClass’s refined core class correctly identifies the optimal core class, which is “library”. In the other example, TELEClass also pinpoints the most accurate core class “bathroom aids safety” while other methods can only find more general or partially relevant classes.
5 Related Work
5.1 Weakly-Supervised Text Classification
Weakly-supervised text classification trains a classifier with limited guidance, aiming to reduce human efforts while maintaining high proficiency. Various sources of weak supervision have been explored, including distant supervision Gabrilovich and Markovitch (2007); Chang et al. (2008); Song and Roth (2014) like knowledge bases, keywords Agichtein and Gravano (2000); Tao et al. (2018); Meng et al. (2018, 2020); Wang et al. (2021a); Zhang et al. (2021a), and heuristic rules Ratner et al. (2016); Badene et al. (2019); Shu et al. (2020). Later, extremely weakly-supervised methods are proposed to solely rely on class names to generate pseudo labels and train classifiers. LOTClass Meng et al. (2020) utilizes MLM-based PLM as a knowledge base for extracting class-indicative keywords. X-Class Wang et al. (2021a) extracts keywords for creating static class representations through clustering. PIEClass Zhang et al. (2023c) employs PLMs’ zero-shot prompting to obtain pseudo labels with noise-robust iterative ensemble training. MEGClass Kargupta et al. (2023) acquires contextualized sentence representations to capture topical information at the document level. WOTClass Wang et al. (2023) ranks class-indicative keywords from generated classes and extracts classes with overlapping keywords.
5.2 Hierarchical Text Classification
A hierarchy provides a systematic top-to-down structure with inherent semantic relations that can assist in text classification. Typical hierarchical text classification can be categorized into two groups: local approaches and global approaches. Local approaches train multiple classifiers for each node or local structures Banerjee et al. (2019); Liu et al. (2005); Wehrmann et al. (2018). Global approaches, learn hierarchy structure into a single classifier through recursive regularization Gopal and Yang (2013), a graph neural network (GNN)-based encoder Peng et al. (2018); Huang et al. (2019); Zhou et al. (2020); Shen et al. (2021), or a joint document label embedding space Chen et al. (2021).
Weak supervision is also studied for hierarchical text classification. WeSHClass Meng et al. (2019) uses a few keywords or example documents per class and pretrains classifiers with pseudo documents followed by self-training. TaxoClass Shen et al. (2021) follows the same setting as ours which uses the sole class name of each class as the only supervision. It identifies core classes for each document using a textual entailment model, which is then used to train a multi-label classifier. Additionally, MATCH Zhang et al. (2021c) and HiMeCat Zhang et al. (2021b) study how to integrate associated metadata into the label hierarchy for document categorization with weak supervision.
5.3 LLMs as Generators and Annotators
Large language models (LLMs) have demonstrated impressive performance in many downstream tasks and are explored to help low-resource settings by synthesizing data as generators or annotators Ding et al. (2023). For data generation, few-shot examples Wang et al. (2021b) or class-conditioned prompts Meng et al. (2022) are explored for LLM generation and the generated data can be used as pseudo training data to further fine-tune a small model as the final classifier Ye et al. (2022). Recently, Yu et al. (2023) proposed an attribute-aware topical text classification method that incorporates ChatGPT to generate topic-dependent attributes and topic-independent attributes to reduce topic ambiguity and increase topic diversity for generation. For data annotation, previous works utilize LLMs for unsupervised annotation Feng et al. (2021), Chain-of-Thought annotation with explanation generation He et al. (2024), and active annotation Zhang et al. (2023a).
6 Conclusion and Future Work
In this paper, we propose a new method, TELEClass, for the minimally-supervised hierarchical text classification task with two major contributions. First, we enrich the input label taxonomy with LLM-generated and corpus-based class-indicative terms for each class, which can serve as additional features to understand the classes and facilitate classification. Second, we explore the utilization of LLMs in the hierarchical text classification in two directions: data annotation and data creation. On two public datasets, TELEClass can outperform existing baseline methods substantially, and we further demonstrate its effectiveness through ablation studies. We also conduct a comparative analysis of performance and cost for zero-shot LLM prompting for the hierarchical text classification task. For future works, first, we plan to generalize TELEClass’s idea into other low-resource text mining tasks with hierarchical label spaces, such as fine-grained entity typing. We also plan to explore how to extend TELEClass into more complicated label spaces, such as an extremely large space with millions of labels.
Limitations
In this paper, we propose TELEClass, a novel hierarchical text classification method that uses the label taxonomy as the only supervision. Despite its effectiveness shown in the experiments, there is still room to improve. First, TELEClass utilizes the general knowledge of pretrained LLM and shows strong performance on two general domain datasets. However, it remains a challenge to train effective text classifiers in private domains, for which LLMs lack relevant knowledge. Second, in this paper, we mainly focus on acquiring high-quality pseudo labeled data while only utilizing the simplest classifier model and objective. It is worth studying how the proposed method can be further improved with more advanced network structure and noise-robust training objectives. Lastly, TELEClass assumes the availability of a large unlabeled corpus for pseudo labeling, whose performance is not fully studied if limited data is available.
Ethics Statement
In this work, we develop a new hierarchical text classification method and experiment on two publicly available datasets for research purposes. Overall, we do not foresee any significant ethical issues from our work. All the datasets are used as intended. We will open-source our code upon publication to assist future studies. Although showing strong language understanding capability, LLMs can sometimes generate misinformation Chen and Shu (2023) or biased contents Fang et al. (2024). In our work, instead of studying the generation ability of LLMs, we only use LLMs to generate pseudo documents for text classifier training. In practice, our method can be combined with any misinformation detection and debiasing methods Mattern et al. (2022); Chen and Shu (2021).
Acknowledgments
Research was supported in part by US DARPA INCAS Program No. HR001121C0165 and KAIROS Program No. FA8750-19-2-1004, National Science Foundation IIS-19-56151, and the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No. 2019897, and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No. 2118329. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of DARPA or the U.S. Government.
Appendix A Pseudo code of TELEClass
Algorithm 1 summarizes the entire TELEClass framework.